OpenStudy (anonymous):

Given f′′(x)=2x+4 and f′(−3)=0 and f(−3)=−1. find f'(x) and f(3) Hint: start with integrating the second derivative function, and use the "fixed point" f′(−3)=0. Then solve for C, giving you the first derivative function. Now integrate this function, and use the second fixed point given to solve the second part of the problem.

OpenStudy (anonymous):

\[f''(x)=2x+4\] \[f'(x)=x^2+4x+C\]

OpenStudy (anonymous):

you know that \[f'(-3)=0\] so you know \[(-3)^2+4\times -3+C=0\] \[9-12+C=0\] \[C=3\] so \[f'(x)=x^2+4x+3\]

OpenStudy (anonymous):

now repeat the process for the next one

OpenStudy (anonymous):

\[f(x)=\frac{1}{3}x^3+2x^2+3x+C\] and put \[f(-3)=-1=\frac{1}{3}(-3)^3+2(-3)^2+3(-3)+C\] and solve for C

OpenStudy (anonymous):

not getting it. I get 5 for C, then what?

OpenStudy (anonymous):

then i guess you are done right?

OpenStudy (anonymous):

\[-9+18-9+C=-1\] \[C=-1\] is what i got, n ot 5

OpenStudy (anonymous):

then i guess you are done right?

OpenStudy (anonymous):

I'm confused. Why solve for f(-3) when it's asking for f(3)?

OpenStudy (anonymous):

ok i did not solve for f(-3)

OpenStudy (anonymous):

it says and f(−3)=−1. and i used that to solve for C. i got C = -1

OpenStudy (anonymous):

so how would I get f(3) then? nothing that I'm doing seems to work. Last thing I tried to input was ((x^3+6x^2)/3)-1

OpenStudy (anonymous):

ok i think we have \[f(x)=\frac{1}{3}x^3+2x^2++3x-1\] yes?

OpenStudy (anonymous):

sorry i mean \[f(x)=\frac{1}{3}x^3+2x^2+3x-1\]

OpenStudy (anonymous):

unless i made a mistake somewhere with the constants. now to find \[f(x)=\frac{1}{3}3^3+2\times 3^2+3\times 3-1\] i get \[9+18+9-1=35\]

OpenStudy (anonymous):

ok, that worked. Now I'm gonna take some time to see where I went wrong. Thanks